OpenStudy (anonymous):

Find f if g(x)=9x+(3/x) and (fog)(x)=81x^2+(9/x^2)

OpenStudy (anonymous):

fog(x) as in fg(x) ?

OpenStudy (loser66):

\[f(x) = x^2 -54\] If you want to know how to get it, make question.

OpenStudy (anonymous):

Is it the same form used in defining chain rule? F(x)=(f o g)(x), F'(x)=f'(g(x))*g'(x). I think this problem needs linear algebra, which however I almost forget. There probably more than one answer.

OpenStudy (anonymous):

But the above guy is actually right, the simple observation gives at least one answer, R(x)=81x^2+(9/x^2)=f(g(x))=g(x)^2-54, so f(x)=x^2-54 is one answer.